• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name kominis

    Log entry time 16:00:49 on October 14,1998

    Entry number 2858

    keyword=13:00 - 16:00 Summary

    14:40 Magnets have settled at 0.666 GeV/c
    Asked for controlled access because there was a network problem
    of Moller data aquisition.
    M. Jones will work at the hadron arm, to change the h+ helicity pulser channel.
    15:45 : Moller work in the hall done.
    Target Polarization NMR Signal 11.5 mV ( approximately 35%)
    16:01 : Work finished.
    X. Jiang checked the hadron arm lead glass #23, signal looks OK.
    High voltage might need to be changed later.